Consider the discrete random variables X and Y with the following joint probability
mass function:

T -1 0 0 Y 0 -1 1 0 1 -1 fxy(x,y) 1/8 1/4 1/4 1/8 1/8 1/8 1 다.

Given that X is not negative, what is the probability that Y is also not negative?
A. 0.5 B. 0.8 C. 0.4 D. 0.25 E. none of the preceding
Answer #1

Given that X is not negative, the probability that Y is also not negative is computed using Bayes theorem here as:

P(Y >= 0 | X >= 0) = P(X >= 0, Y >= 0) / P(X >= 0)

= (0.25 + 0.125) / (0.25 + 0.25 + 0.125 + 0.125)

= 0.375 / 0.75

= 0.5

Therefore 0.5 is the required probability here. ( A is the correct answer here )